# Tracing quantum correlations back to collective interferences

### v1

@@ Abstract
In this paper, we investigate the possibility of explaining nonclassical correlations between two quantum systems in terms of quantum interferences between collective states of the two systems. We achieve this by mapping the relations between different measurement contexts in the product Hilbert space of a pair of two-level systems onto an analogous sequence of interferences between paths in a single-particle interferometer. The paradoxical relations between different measurement outcomes can then be traced to the distribution of probability currents in the interferometer. We show that the relation between probability currents and correlations can be represented by continuous conditional (quasi)probability currents through the interferometer, given by weak values; the violation of the noncontextual assumption is expressed by negative conditional currents in some of the paths. Since negative conditional currents correspond to the assignment of negative conditional probabilities to measurements results in different measurement contexts, the necessity of such negative probability currents represents a failure of noncontextual local realism. Our results help to explain the meaning of nonlocal correlations in quantum mechanics, and support Feynman’s claim that interference is the origin of all quantum phenomena.
@@ I Introduction
In order to examine the aspects of the quantum formalism which are behind the paradoxical behaviours associated with nonlocality and contextuality, two of the authors previously analysed the quantitative relations between the different measurement contexts defined by the Hilbert space formalism [19, 20, 21]. Ref. [20] showed that quantum contextuality and nonlocality can both be demonstrated through the same quantitative relations between measurement contexts, for contexts defined by a set of orthogonal bases in a single-particle three-dimensional Hilbert space, and a set of local-state orthogonal bases in a product Hilbert space describing a pair of two-level systems, respectively. In parallel, Ref. [21] introduced a three-path interferometer, in which the five measurement contexts required for the demonstration of quantum contextuality are physically implemented by a sequence of beam splitters. Each of the paths represents a possible measurement outcome obtained in one of the measurement contexts, directly relating the problem of contextuality to the problem of attempting to trace the path of a photon through the interferometer. The abstract problem of relating different measurement contexts to each other using their Hilbert space representation can thus be illustrated by the propagation of a photon through an interferometer. Ref. [20] also showed that the relation between two entangled qubits can be expressed in a three-dimensional subspace of above-mentioned four-dimensional product space, when collective measurements of the two systems are included in the description. This equivalence means that we can map the relations between measurement contexts for qubit pairs onto the three-dimensional Hilbert space described by the three-path interferometer. It is then possible to explore the quantum correlations between two qubits using concepts that were previously limited to the investigation of single-particle contextuality.
@@
Feynman famously claimed that interference is the origin of all quantum phenomena [22]. Ref. [21], which illustrates contextuality through single particle interferences, provides support for this view. However, it is not immediately obvious how quantum interferences relate to nonclassical correlations between separate quantum systems [23, 24, 25]. Taken together, Ref. [20] (relating contextuality and nonlocality) and Ref. [21] (relating contextuality to interference) suggest a potential route to clarify the relation between nonlocal quantum correlations, and quantum interference between the amplitudes associated with different measurement outcomes. The specific problem that makes it difficult to identify quantum interference effects in both Bell inequalities and Hardy’s paradox originates from the assumption of local realism [12], which requires that the alternative measurement contexts all correspond to local measurements. Here, we instead focus on the role of quantum interference in relating contexts to one another, regardless of whether these contexts are locally defined or not. This allows us to take a “shortcut” through Hilbert space by using the outcomes of collective measurements to relate local measurement outcomes to each other. We can then explain the physics of nonclassical correlations in terms of the collective interference between states representing the joint properties of the two quantum systems.
@@
In Section V, we relate the distribution of probability currents to the quasiprobabilities given by the Kirkwood-Dirac distribution [26, 27, 28], allowing us to identify the effects of interferences directly with nonclassical correlations between different measurement contexts. Paradoxical statistics are described by negative weak values, corresponding to negative joint probabilities in the Kirkwood-Dirac distribution [29, 30, 31, 32]. We show that the paradoxical correlations associated with the entanglement of the two qubits can be traced back to a negative conditional current at a collective measurement outcome. It is possible to explain the role of these collective measurement outcomes by considering the local states to which they are connected by the beam splitters in the interferometer. Each collective measurement outcome can be identified with a set of conditional statements about the two qubits that must be valid if that outcome is detected. The appearance of negative conditional currents shows that it is impossible to assign simultaneous reality to both the collective statements and the local statements represented by the different measurement contexts. Quantum interferences thus trace the origin of quantum correlations characteristic of entangled states back to the fundamental dependence of the reality of the measurement outcomes on the context established by the application of a specific measurement. Finally, in Section VI, we discuss the implications of our work, and summarise the paper.
@@ II Illustration of collective interference using an analogous interferometer
On the other hand, the output state $\left|N_{1}\right\rangle$ is not symmetric in the two systems, and the swap operation changes it into a different state, $\left|N_{2}\right\rangle$. We can set the reflectivity $R_{\mathrm{f}}$ of the next beam splitter to $R_{\mathrm{f}}=1/4$, so that the interference between $\left|N_{1}\right\rangle$ and $\left|b,0\right\rangle$ corresponds to the swap operation acting on the two systems,
@@ III Relation between local and collective measurement outcomes
As shown in [21], the interferometer given in Fig. 1 can be used to demonstrate quantum contextuality, by relating different measurement contexts to each other sequentially. Here, we apply this tool to the problem of nonclassical correlations between two quantum systems. We follow the analogy between contextuality and nonlocality previously established in [20]. As shown in that work, one could extend the interferometer to fully represent a version of Hardy’s paradox. However, we instead intend to show that the three-path interferometer analogy is sufficient to demonstrate collective quantum effects in the two systems, based on the non-separability of the input state. We will do this by investigating how interference describes the relation of the central collective outcome $\left|f_{\text{NL}}\right\rangle$ to local outcomes belonging to different contexts.
@@
This inequality quantifies the relation between the local and collective measurement outcomes required by the “either/or” interpretation of quantum interference. Note that this is a typical noncontextual inequality, of the same form as given in [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 19, 44, 45]. It is well-known that quantum mechanics can violate such inequalities. Fig. 2 gives an idea of how this violation can be achieved. Quantum interferences should increase the probability of $\left|f_{\text{NL}}\right\rangle$ beyond the value of the sum of the probabilities of $\left|a,0\right\rangle$ and $\left|0,a\right\rangle$. Since the superposition of $\left|0,1\right\rangle$ and $\left|1,0\right\rangle$ can provide such an enhancement while limiting the probabilities of $\left|a,0\right\rangle$ and $\left|0,a\right\rangle$ to only half the probabilities of $\left|0,1\right\rangle$ and $\left|1,0\right\rangle$ respectively, the ideal state for a violation of the inequality should be
@@ IV Collective interference and quantum correlations
The interference effect redistributes probability between $P(f_{\text{NL}})$, and $P(N_{1})$ and $P(N_{2})$. The noncontextual conditions given by Statements 1 and 2 can be satisfied if the current through $\left|N_{1}\right\rangle$ and $\left|N_{2}\right\rangle$ is sufficiently high, to establish a connection between $\left|b,0\right\rangle$ and $\left|1,0\right\rangle$, and between $\left|0,1\right\rangle$ and $\left|0,b\right\rangle$, that does not involve either $\left|a,0\right\rangle$ or $\left|0,a\right\rangle$. If collective interference increases $P(f_{\mathrm{NL}})$ beyond the threshold of 1/2, it is difficult to identify a continuous probability current through the interferometer representing the collective quantum interference of the two systems. The extremal case associated with the input state $\left|\Phi_{\mathrm{max}}\right\rangle$ (where $\eta=1$) is shown in Fig. 3. For this case, the probabilities of $\left|N_{1}\right\rangle$ and $\left|N_{2}\right\rangle$ are each only 1/12. In the figure, we can see it looks as though a current from $\left|1,0\right\rangle$ to $\left|1,0\right\rangle$ should pass through either $\left|a,0\right\rangle$ or $\left|N_{2}\right\rangle$—however, the sums of the probabilities $P(a,0)$ and $P(N_{2})$ are not sufficient to account for the probability of obtaining output $\left|1,0\right\rangle$.

### v2

@@ Abstract
In this paper, we investigate the possibility of explaining nonclassical correlations between two quantum systems in terms of quantum interferences between collective states of the two systems. We achieve this by mapping the relations between different measurement contexts in the product Hilbert space of a pair of two-level systems onto an analogous sequence of interferences between paths in a single-particle interferometer. The relations between different measurement outcomes are then traced to the distribution of probability currents in the interferometer, where paradoxical relations between the outcomes are identified with currents connecting two states that are orthogonal and should therefore exclude each other. We show that the relation between probability currents and correlations can be represented by continuous conditional (quasi)probability currents through the interferometer, given by weak values; the violation of the noncontextual assumption is expressed by negative conditional currents in some of the paths. Since negative conditional currents correspond to the assignment of negative conditional probabilities to measurements results in different measurement contexts, the necessity of such negative probability currents represents a failure of noncontextual local realism. Our results help to explain the meaning of nonlocal correlations in quantum mechanics, and support Feynman’s claim that interference is the origin of all quantum phenomena.
@@ I Introduction
In order to examine the aspects of the quantum formalism which are behind the paradoxical behaviours associated with nonlocality and contextuality, two of the authors previously analysed the quantitative relations between the different measurement contexts defined by the Hilbert space formalism [19, 20, 21]. Ref. [20] showed that quantitative arguments about contextuality based on Hilbert space inner products can be used equally to analyse a three-dimensional Hilbert space of a single system or the nonlocality of two entangled two-level systems. In parallel, Ref. [21] introduced a three-path interferometer, in which the five measurement contexts required for the demonstration of quantum contextuality in a three-dimensional Hilbert space are physically implemented by a sequence of beam splitters. Each of the paths represents a possible measurement outcome obtained in one of the measurement contexts represented by a quantum state, directly relating the problem of contextuality to the problem of attempting to trace the path of a photon through the interferometer. Ref. [20] also showed that the nonclassical correlations between two entangled qubits can be expressed in a three-dimensional Hilbert space when collective measurements of the two systems are included in the description. This means that we can map the relations between measurement contexts for qubit pairs onto the path of photon through the three-path interferometer. It is then possible to explore the quantum correlations between two qubits using concepts that were previously limited to the investigation of single-particle contextuality.
@@
Feynman famously claimed that interference is the origin of all quantum phenomena [22]. Ref. [21], which illustrates contextuality through single particle interferences, provides support for this view. However, it is not immediately obvious how quantum interferences relate to nonclassical correlations between separate quantum systems [23, 24, 25, 26]. Taken together, Ref. [20] (relating contextuality and nonlocality) and Ref. [21] (relating contextuality to interference) suggest a potential route to clarify the relation between nonlocal quantum correlations, and quantum interference between the amplitudes associated with different measurement outcomes. The specific problem that makes it difficult to identify quantum interference effects in both Bell inequalities and Hardy’s paradox originates from the assumption of local realism [12], which requires that the alternative measurement contexts all correspond to local measurements. Focusing on the quantum mechanical description of the relation between two systems, it is important to analyze the meaning of collective measurement outcomes through their expression as a superposition of local states. We can then take a “shortcut” through Hilbert space by using the outcomes of collective measurements to relate local measurement outcomes to each other. The physics of nonclassical correlations between two systems can thus be described by the different ways in which the same collective state can be expressed by different superpositions of local states.
@@
In Section V, we relate the distribution of probability currents to the quasiprobabilities given by the Kirkwood-Dirac distribution [27, 28, 29], allowing us to identify the effects of interferences directly with nonclassical correlations between different measurement contexts. Paradoxical statistics are described by negative weak values, corresponding to negative joint probabilities in the Kirkwood-Dirac distribution [30, 31, 32, 33, 34]. We show that the paradoxical correlations associated with the entanglement of the two qubits can be traced back to a negative conditional current at a collective measurement outcome. It is possible to explain the role of these collective measurement outcomes by considering the local states to which they are connected by the beam splitters in the interferometer. Each collective measurement outcome can be identified with a set of conditional statements about the two qubits that must be valid if that outcome is detected. The appearance of negative conditional currents shows that it is impossible to assign simultaneous reality to both the collective statements and the local statements represented by the different measurement contexts. Quantum interferences thus trace the origin of quantum correlations characteristic of entangled states back to the fundamental dependence of the reality of the measurement outcomes on the context established by the application of a specific measurement. Finally, in Section VI, we discuss the implications of our work, and summarise the paper.
@@ II Illustration of collective interference using an analogous interferometer
On the other hand, the output state $\left|N_{1}\right\rangle$ is not symmetric in the two systems, and the swap operation changes it into a different state $\left|N_{2}\right\rangle$. Using the reflectivities given above, their representations in the input basis read,
$\begin{split}&\left|N_{1}\right\rangle=\frac{1}{\sqrt{6}}(\left|0,0\right\rangle+2\left|0,1\right\rangle-\left|1,0\right\rangle)\\ &\left|N_{2}\right\rangle=\frac{1}{\sqrt{6}}(\left|0,0\right\rangle-\left|0,1\right\rangle+2\left|1,0\right\rangle).\end{split}$ (3)
The inner product of these two states is $-1/2$, indicating that the reflectivity $R_{\mathrm{f}}$ of the next beam splitter should be $1/4$, so that the interference between $\left|N_{1}\right\rangle$ and $\left|b,0\right\rangle$ corresponds to the swap operation acting on the two systems,
@@ III Relation between local and collective measurement outcomes
As shown in [21], the interferometer given in Fig. 1 can be used to demonstrate quantum contextuality, by relating different measurement contexts to each other sequentially. Here, we apply this tool to the problem of nonclassical correlations between two quantum systems. We follow the analogy between contextuality and nonlocality previously established in [21, 20]. As shown in that work, one could extend the interferometer to fully represent a version of Hardy’s paradox. However, we instead intend to show that the three-path interferometer analogy is sufficient to demonstrate collective quantum effects in the two systems, based on the non-separability of the input state. We will do this by investigating how interference describes the relation of the central collective outcome $\left|f_{\text{NL}}\right\rangle$ to local outcomes belonging to different contexts.
@@
Note that this is a typical noncontextual inequality, of the same form as given in [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 19, 46, 47]. This inequality quantifies the limit of the “either/or” interpretation associated with the detection of single particles. It has been pointed out that its violation can be explained by the intensity distributions of classical wave interferences [48]. As explained in [21], this is a consequence of the joint presence of all amplitudes and intensities of a classical wave, which makes it impossible to identify the origin of different energy terms in the output. Since a single particle always arrives at only one output port, it is natural to assume that it also entered at only one port. The violation of the inequality indicates that this assumption might be problematic.
It is central that quantum interference can modify the contributions of the different outcomes in a manner that contradicts “either/or” logic, resulting in a violation of the inequality. Fig. 2 gives an idea of how this violation can be achieved. Quantum interferences should increase the probability of $\left|f_{\text{NL}}\right\rangle$ beyond the value of the sum of the probabilities of $\left|a,0\right\rangle$ and $\left|0,a\right\rangle$. Since the superposition of $\left|0,1\right\rangle$ and $\left|1,0\right\rangle$ can provide such an enhancement while limiting the probabilities of $\left|a,0\right\rangle$ and $\left|0,a\right\rangle$ to only half the probabilities of $\left|0,1\right\rangle$ and $\left|1,0\right\rangle$ respectively, the ideal state for a violation of the inequality should be
@@ IV Collective interference and quantum correlations
The interference effect redistributes probability between $P(f_{\text{NL}})$, and $P(N_{1})$ and $P(N_{2})$. The noncontextual conditions given by Statements 1 and 2 can be satisfied if the current through $\left|N_{1}\right\rangle$ and $\left|N_{2}\right\rangle$ is sufficiently high, to establish a connection between $\left|b,0\right\rangle$ and $\left|1,0\right\rangle$, and between $\left|0,1\right\rangle$ and $\left|0,b\right\rangle$, that does not involve either $\left|a,0\right\rangle$ or $\left|0,a\right\rangle$. If collective interference increases $P(f_{\mathrm{NL}})$ beyond the threshold of 1/2, it is difficult to identify a continuous probability current through the interferometer representing the collective quantum interference of the two systems. The extremal case associated with the input state $\left|\Phi_{\mathrm{max}}\right\rangle$ (where $\eta=1$) is shown in Fig. 3. For this case, the probabilities of $\left|N_{1}\right\rangle$ and $\left|N_{2}\right\rangle$ are each only 1/12. In the figure, we can see it looks as though a current from $\left|1,0\right\rangle$ to $\left|1,0\right\rangle$ should pass through either $\left|a,0\right\rangle$ or $\left|N_{2}\right\rangle$—however, the sums of the probabilities $P(a,0)$ and $P(N_{2})$ are not sufficient to account for the probability of obtaining output $\left|1,0\right\rangle$. This suggests that part of the probability current in the output port $\left|1,0\right\rangle$ should be transferred from the input port $\left|0,1\right\rangle$.